OpenStudy (anonymous):

Check my answer? Find S12 for the series 2 + 5 + 8 +... I think it's 35

jimthompson5910 (jim_thompson5910):

S12 means "the 12th partial sum". It's a lot bigger than 35

OpenStudy (anonymous):

Can you explain how to do it please?

jimthompson5910 (jim_thompson5910):

what is the nth term equal to

OpenStudy (anonymous):

2?

jimthompson5910 (jim_thompson5910):

the nth term will be some expression with n in it

OpenStudy (anonymous):

okay is it 2n then? I'm sorry I'm confused on the section I'm doing.

jimthompson5910 (jim_thompson5910):

first term is a1 = 2 common difference is d = 3 you plug those things into an = a1 + d(n-1) to get the nth term

OpenStudy (anonymous):

okay I got n=1? I feel like I somehow did something wrong in that equation lol

jimthompson5910 (jim_thompson5910):

\[\Large a_{n} = a_1 + d(n-1)\] \[\Large a_{n} = 2 + 3(n-1)\] \[\Large a_{n} = ???\]

OpenStudy (anonymous):

okay thank you I understand the equation now

jimthompson5910 (jim_thompson5910):

so that simplifies to what?